(a) A random variable X has the following probability function

.x : -2 -1 0 1 2 3

p(x): 0.1 k 0·2 2k 0·3 k

(i) Find the value of k. and calculate mean and variance.

(ii) construct the cumulative distribution function c.d.f F(x) and draw its graph.


1
Expert's answer
2020-10-12T18:57:08-0400

Solution to i

"\sum p(x)=1"

Thus;

"0.1+k+0.2+2k+0.3+k=1""0.6+4k=1 \therefore 4k=0.4""k=0.1"

Mean, "\bar{x}"

"\bar{x}=\sum xp(x)"

"=-2(0.1)-1(0.1)+0(0.2)+1(0.2)+2(0.3)+3(0.1)""=0.8"

variance

"var= \sum x^2p(x)- {\bar{x}}^2"

"\sum x^2p(x)" "=4(0.1)+1(0.1)+0(0.2)+1(0.2)+4(0.3)+9(0.1)"

"=2.8"


"var=2.8-({0.8^2})=2.16"

Solution to ii

Cumulative Distribution Function



Cumulative Distribution Graph